The table below represents the costs for a computer company for a week (sorry, they’re not...

  1. The table below represents the costs for a computer company for a week (sorry, they’re not very realistic).

Quantity

Fixed Costs

Variable Costs

Total Costs

Average Total Cost

Average Variable Costs

Marginal Cost

1

$1000

$600

                 

2

$1000

$1100

3

$1000

$1500

4

$1000

$1800

5

$1000

$2200

6

$1000

$2700

7

$1000

$3400

8

$1000

$4500

  1. Complete the missing information in the table.
  2. At what price would the company shut-down in the short run?
  3. Suppose this economy is operating in a perfectly competitive market. If the market price of a computer is $750, how many computers will the firm want to produce and sell to maximize profit? How much profit will the firm get?
  4. Assume the cost structure of this firm is typical for this industry. Do you expect firms to enter or exit the market? Why or why not? How will this affect profits of this firm?
  5. Assume that this is a constant cost industry. What do you expect the price for a computer to be in the long run? Explain how you know.
Answer #1
Q FC VC TC ATC AVC MC
1 1000 600 1600 1600 600 600
2 1000 1100 2100 1050 550 500
3 1000 1500 2500 833.3333 500 400
4 1000 1800 2800 700 450 300
5 1000 2200 3200 640 440 400
6 1000 2700 3700 616.6667 450 500
7 1000 3400 4400 628.5714 485.7143 700
8 1000 4500 5500 687.5 562.5 1100

a) TC=FC+VC

ATC=TC/Q

AVC=VC/Q

MC=change in TC/change in Q

b) The company will shut down at any price below its minimum AVC = 440

c) P=750

The firm will set P=MC for profit maximization so it will produce Q=7 units

profits = TR-TC = (750*7)-4400 = 850

d) More firms will enter the market in the long run as the industry is earning positive profits.

As more firms enter, the supply increases which decreases the price until it is equal to the minimum ATC so the profits becomes zero in the long run and the firms only earn normal profits.